OpenStudy (anonymous):

you roll 2 dice, what is the probability that at least one of them is a 6?

OpenStudy (anonymous):

1/6

hartnn (hartnn):

P(atleast 1 6) =P{(6 on 1st dice AND not 6 on other)OR(6 on 2nd dice AND not 6 on first)OR(6 on 1st dice AND 6 on other) =1/6 * 5/6 + 1/6* 5/6 + 1/6*1/6 =??

OpenStudy (anonymous):

The probability of an event \(E\in S\) NOT happening is: \(1-P(E)\). We use this to account for the probability of both die *not* rolling a 6, which is: \[ P(S\setminus E_6)=1-\frac{1}{6}=\frac{5}{6} \]Thus, we make use of the probability to find: \[ P(S\setminus E_6)\cdot P(S\setminus E_6)=\left(\frac{5}{6}\right)^2=\frac{25}{36} \]Finally, we compute the true probability of the event of rolling both die as a six \(E_{6,6}\) happening, to receive: \[ P(E_{6,6})=1-\frac{25}{36}=\frac{11}{36} \]Thus we are done.

OpenStudy (anonymous):

"rolling both die as a six"->"rolling both die and receiving a six"

OpenStudy (zarkon):

I don't think you want to write \(E\in S\)

OpenStudy (anonymous):

I meant to say, for some event-space \(S\) and an event \(E\in S\).

OpenStudy (zarkon):

\(E\subseteq S\)

OpenStudy (zarkon):

events are sets

OpenStudy (anonymous):

Yes, thank you, my bad. I always abuse notation, just forget.

OpenStudy (zarkon):

otherwise...nice solution.

OpenStudy (anonymous):

Grazie

OpenStudy (anonymous):

thanks very much everyone, HUGE help!!!!! =)